# CARI4D: Category Agnostic 4D Reconstruction of Human-Object Interaction

### v1

@@ 3 Method
Specifically, given a sequence of $N$ RGB images $\{\mathbf{I}_{i}\}_{i=1}^{N}$ where a person interacts with an object, our goal is to reconstruct the object mesh $\mathbf{O}$ with corresponding per-frame 6DoF poses $\{\mathcal{O}_{i}\}_{i=1}^{N}$, here $\mathcal{O}_{i}=(\mathbf{R}^{o},\mathbf{t}^{o})$ consists of rotation $\mathbf{R}^{o}\in SO(3)$ and translation $\mathbf{t}^{o}\in\mathbb{R}^{3}$. For the human, we use the SMPL-H body model [25, 35] and estimate per-frame parameters $\{\mathcal{H}_{i}\}_{i=1}^{N}$, where $\mathcal{H}_{i}=(\boldsymbol{\theta},\beta,\mathbf{t}^{h})$ consists of the SMPL body pose $\boldsymbol{\theta}$, shape $\beta$ and global translation $\mathbf{t}^{h}$ parameters. Note that we reconstruct the human and object at metric scale with consistent global translations as this is important for downstream applications.
@@ 3.2 Human and Object Pose Initialization
Object pose hypothesis selection. FoundationPose internally produces $K$ pose candidates and their quality scores per frame, selecting the pose with the highest score as the output. In our considered more challenging setup, its top-1 prediction often fails, though the correct pose typically lies within the $K$ candidates. Our key idea is to dynamically select the optimal pose per frame by leveraging visual cues and temporal consistency. Specifically, given object pose $\hat{\mathcal{O}}_{i-1}$ at previous frame i1i-1, our goal is to select the best pose from a list of pose hypothesis $\{\hat{\mathcal{O}}_{i}^{j}\}_{j=1}^{K}$ (ranked from best to worst by FoundationPose) for current frame $i$. We consider two criteria to filter out wrong poses: a). mask IoU and b). temporal smoothness.
@@ 3.4 Contact-based Joint Optimization
Our feedforward prediction from CoCoNet improves the relative pose between human and object, yet it cannot guarantee the contacts are satisfied and predictions are image aligned. We further improve the contact coherency and motion smoothness via contact aware joint optimization, guided by the contact predictions from Sec. 3.3. Specifically, we optimize the human and object pose parameters $\{\mathcal{H}_{i},\mathcal{O}_{i}\}_{i=1}^{N}$ by minimizing the contact distances while satisfying visual constraints from input. The objective function is defined as:
@@ 6.2 Joint Optimization
The 2D object mask loss $L_{m}$ is occlusion aware and defined as: $L_{m}=\sum(M-I\circ S)^{2}$, where $M$ is the input object mask, $I$ is a non-occlusion indicator (1 if pixel belongs only to this object, 0 else) and $S$ is the rendered object silhouette. The non-occlusion indicator is derived from human and object masks. The human and object masks may have overlap regions due to imperfect segmentation. To address this, pixels that belong only to object mask are assigned value 1 for the non-occlusion indicator. This avoids computing loss on regions where the object is occluded, see illustration in PHOSA [73].

### v2

@@ 3 Method
Specifically, given a sequence of NN RGB images $\{\mathbf{I}_{i}\}_{i=1}^{N}$ where a person interacts with an object, our goal is to reconstruct the object mesh $\mathbf{O}$ with corresponding per-frame 6DoF poses $\{\mathcal{O}_{i}\}_{i=1}^{N}$, here $\mathcal{O}_{i}=(\mathbf{R}^{o},\mathbf{t}^{o})$ consists of rotation $\mathbf{R}^{o}\in SO(3)$ and translation $\mathbf{t}^{o}\in\mathbb{R}^{3}$. For the human, we use the SMPL-H body model [25, 35] and estimate per-frame parameters $\{\mathcal{H}_{i}\}_{i=1}^{N}$, where $\mathcal{H}_{i}=(\boldsymbol{\theta},\beta,\mathbf{t}^{h})$ consists of the SMPL body pose $\boldsymbol{\theta}$, shape $\beta$ and global translation $\mathbf{t}^{h}$ parameters. Note that we reconstruct the human and object at metric scale with consistent global translations as this is important for downstream applications.
@@ 3.2 Human and Object Pose Initialization
Object pose hypothesis selection. FoundationPose internally produces $K$ pose candidates and their quality scores per frame, selecting the pose with the highest score as the output. In our considered more challenging setup, its top-1 prediction often fails, though the correct pose typically lies within the $K$ candidates. Our key idea is to dynamically select the optimal pose per frame by leveraging visual cues and temporal consistency. Specifically, given object pose $\hat{\mathcal{O}}_{i-1}$ at previous frame i1i-1, our goal is to select the best pose from a list of pose hypothesis $\{\hat{\mathcal{O}}_{i}^{j}\}_{j=1}^{K}$ (ranked from best to worst by FoundationPose) for current frame ii. We consider two criteria to filter out wrong poses: a). mask IoU and b). temporal smoothness.
@@ 3.4 Contact-based Joint Optimization
Our feedforward prediction from CoCoNet improves the relative pose between human and object, yet it cannot guarantee the contacts are satisfied and predictions are image aligned. We further improve the contact coherency and motion smoothness via contact aware joint optimization, guided by the contact predictions from Sec. $3.$ Specifically, we optimize the human and object pose parameters $\{\mathcal{H}_{i},\mathcal{O}_{i}\}_{i=1}^{N}$ by minimizing the contact distances while satisfying visual constraints from input. The objective function is defined as:
@@ 6.2 Joint Optimization
The 2D object mask loss $L_{m}$ is occlusion aware and defined as: $L_{m}=\sum(M-I\circ S)^{2}$, where $M$ is the input object mask, II is a non-occlusion indicator (1 if pixel belongs only to this object, 0 else) and $S$ is the rendered object silhouette. The non-occlusion indicator is derived from human and object masks. The human and object masks may have overlap regions due to imperfect segmentation. To address this, pixels that belong only to object mask are assigned value 1 for the non-occlusion indicator. This avoids computing loss on regions where the object is occluded, see illustration in PHOSA [73].